### v1

# Monotone Sequences of Metric Spaces with Compact Limits

Abstract.
In this paper, we consider a fixed metric space (possibly an oriented Riemannian manifold with boundary) with an increasing sequence of distance functions and a uniform upper bound on diameter. When the metric space endowed with the pointwise limit of these distances is compact, then there is uniform and Gromov-Hausdorff (GH) convergence to this limit. When the metric space also has an integral current structure of uniformly bounded total mass (as is true for an oriented Riemannian manifold with boundary that has a uniform bound on total volume), we prove volume preserving intrinsic flat convergence to a subset of the GH limit whose closure is the whole GH limit. We provide a review of all notions and have a list of open questions at the end. Dedicated to Xiaochun Rong.
Key words and phrases:
Metric Spaces, Riemannian, Gromov-Hausdorff, Intrinsic Flat
R. Perales’ research was partially funded by NSF DMS 1612049 on Geometric Compactness Theorems.
C. Sormani’s research was partially funded by NSF DMS 1612049 on Geometric Compactness Theorems and a PSC-CUNY grant.
1. Introduction
Our goal is to teach the notions of Gromov-Hausdorff [GH] and Sormani-Wenger intrinsic flat [SWIF] convergence while proving a new theorem that has no assumptions on curvature. The notion of GH distance between metric spaces was first introduced by Edwards in [Edw75] and deeply explored by Gromov in [Gro07] and [Gro81]. See Rong’s 2010 survey [Ron10] for many applications of GH convergence to sequences of Riemannian manifolds with curvature bounds.
The notion of SWIF distance between integral current spaces was introduced by Sormani and Wenger in [SW11] applying the theory of Ambrosio-Kirchheim in [AK00] and work of Wenger in [Wen07]. Volume preserving intrinsic flat ($\mathcal{VF}$) convergence was introduced by Portegies in [Por15]. See Sormani’s survey [Sor17] for many applications of SWIF convergence to sequences of Riemannian manifolds with lower bounds on their scalar curvature. See also papers by Allen, Bryden, Huang, Jauregui, Lakzian, Lee, Perales, Portegies, Sormani, and Wenger which prove SWIF and $\mathcal{VF}$ convergence theorems and present counter examples [SW11][Wen11][LS13] [AS19][AS20] [HLS17][APS24][AP22] [AB24][JL21][JPP22].
The work in this paper is inspired by a theorem in the appendix of [HLS17] by Huang-Lee-Sormani (which applies to sequences with biLipschitz bounds on their distances) and a theorem within [APS24] by Allen-Perales-Sormani (which assumes only volume converging from above and distance from below but requires the limit space to be a compact smooth oriented Riemannian manifold). Here we will only assume the limit space is a compact metric space but we add the hypothesis that the distances are monotone increasing. Our result will be applied by Sormani-Tian in [ST24] to prove SWIF and GH convergence to the extreme limits constructed by Sormani-Tian-Wang in [STW24].
Theorem 1.1.
Given a compact Riemannian manifold, $(M^{m},g_{0})$, possibly with boundary, with a monotone increasing sequence of Riemannian metric tensors $g_{j}$ such that
(1.1) $g_{j}(V,V)\geq g_{j-1}(V,V)\qquad\forall V\in TM$
with uniform bounded diameter,
(1.2) $\operatorname{diam}_{g_{j}}(M)\leq D_{0}.$
Then the induced length distance functions $d_{j}:M\times M\to[0,D_{0}]$ are monotone increasing and converge pointwise to a distance function, $d_{\infty}:M\times M\to[0,D_{0}]$ so that $(M,d_{\infty})$ is a metric space.
If the metric space $(M,d_{\infty})$ is a compact metric space, then $d_{j}\to d_{\infty}$ uniformly and
(1.3) $(M,d_{j})\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}(M,d_{\infty}).$
If, in addition, $M$ is an oriented Riemannian manifold with uniformly bounded total volume,
(1.4) $\operatorname{vol}_{j}(M)\leq V_{0}\textrm{ and }\operatorname{vol}_{j}(\partial M)\leq A_{0}$
then we have volume preserving intrinsic flat convergence
(1.5) $(M,d_{j},[[M]])\stackrel{{\scriptstyle\mathcal{VF}}}{{\longrightarrow}}(M_{\infty},d_{\infty},T_{\infty}).$
where $T_{\infty}$ is an integral current on $(M,d_{\infty})$ such that $T_{\infty}=[[M]]$ viewed as an integral current on $(M,d_{j})$ and
(1.6) $M_{\infty}=\operatorname{set}_{d_{\infty}}(T_{\infty})\subset M\textrm{ with closure }\overline{M}_{\infty}=M.$
Note that Sormani-Wenger already proved that any GH converging sequence satisfying (1.4) has a subsequence converging to a SWIF limit that is a possibly empty subset of the GH limit in [SW11]. In the Riemannian theorem above we use the monotonicity to force the whole sequence to converge and to force the closure of the SWIF limit to agree with the GH limit. We will see in Example 4.14 within that an increasing sequence can converge to a manifold with a cusp singularity, and since a cusp singularity is not included in the definition of an integral current space, this example has $M_{\infty}=\operatorname{set}(T_{\infty})\neq M$.
We will prove this Riemannian theorem in stages as consequences of more general results concerning metric spaces and integral current spaces. We will review the various notions of convergence right before applying them, so that this article may be easily read by a novice. Note that GH, SWIF, and $\mathcal{VF}$ limit spaces are defined using distance functions and do not necessarily agree with the metric completions of limits spaces found by taking smooth convergence of Riemannian metric tensors away from singular sets (see work of Allen-Sormani [AS19]).
In Section 2 we prove a few basic lemmas about metric spaces with monotone increasing sequences of distance functions including a review of the Riemannian distance functions. In Lemma 2.5 we prove pointwise convergence of the monotone increasing sequence of distance functions with a uniform upper bound on diameter. In Example 2.6 we see that we need not have uniform convergence when the limit is noncompact.
In Section 3 we review the definition of Gromov-Hausdorff convergence (Definition 3.1). We then prove the GH and uniform convergence part of Theorem 1.1. In fact we prove a more general theorem concerning a metric space with a monotone increasing sequence of distance functions that converge pointwise to a compact limit in Theorem 3.2. We show the sequence converges uniformly and in the GH sense. We prove this theorem using different techniques than would have been used by Gromov in order to prepare for the SWIF convergence part of the paper. In particular, we construct a common metric space for the converging sequence that has special properties in Proposition 3.9 and Lemma 3.10. In Remark 3.8 we explain that Example 2.6 does not have a Gromov-Hausdorff limit.
In Section 4 we prove the SWIF convergence part of Theorem 1.1. In fact we prove a more general SWIF convergence theorem concerning the convergence of an integral current space with increasing distance functions [Theorem 4.16]. We begin by reviewing De Giorgi’s notion of a Lipschitz tuple [DG95], the notion of a Lipschitz chart, and Ambrosio-Kirchheim’s notion of an integral current and weak convergence [AK00], proving lemmas about how these objects behave for monotone increasing sequences of distances. We review Sormani-Wenger’s definition of an integral current space, including $(M,d,[[M]])$, and their definition of intrinsic flat distance. We then state Theorem 4.16 and apply it to complete the proof of Theorem 1.1 in Section 4.8. Finally we prove Theorem 4.16, first finding a converging subsequence and a limit current structure in Proposition 4.17 and completing the proof in Section 4.10.
In Section 5 we state open problems.
The authors would like to thank Wenchuan Tian and Changliang Wang for inspiring discussions of examples.
2. Pointwise Convergence of Monotone Increasing Distance Functions
Definition 2.1.
A metric space, $(X,d)$, is a collection of points, $X$, paired with a distance function $d:X\times X\to[0,\infty)$ which is definite,
(2.1) $d(x,y)=0\,\,\iff\,\,x=y,$
symmetric,
(2.2) $d(x,y)=d(y,x)\quad\forall x,y\in X,$
and satisfies the triangle inequality,
(2.3) $d(x,y)\leq d(x,z)+d(z,y)\qquad\forall x,y,z\in X.$
Here we consider metric spaces of finite diameter,
(2.4) $\operatorname{diam}_{d}(X)=\sup\{d(x,y):\,x,y\in X\}<\infty.$
We also need the following definition:
Definition 2.2.
We say that $d_{j}:X\times X\to[0,\infty)$ is a monotone increasing sequence of distance functions on $X$, if for all $j\in{\mathbb{N}}$ one has that $(X,d_{j})$ is a metric space and
(2.5) $d_{j+1}(x,y)\geq d_{j}(x,y)\quad\forall x,y\in X.$
2.1. Riemannian Distances
Recall that a Riemannian manifold, $(M,g)$, has a natural metric space, $(M,d_{g})$, defined using the induced Riemannian distance as follows. The lengths of piecewise smooth curves, $C:[0,1]\to M$, are defined by
[…]
and the Riemannian distance between points is defined by
(2.7) $d_{g}(x,y)=\inf\{L_{g}(C)\}$
where the infimum is taken over all piecewise smooth curves, $C:[0,1]\to M$, such that $C(0)=x$ and $C(1)=y$. The diameter of $(M,d_{g})$ is then defined as
(2.8) $\operatorname{diam}_{g}(M)=\operatorname{diam}_{d_{g}}(M)=\sup\{d_{g}(x,y):\,x,y\in M\}.$
Lemma 2.3.
If $M$ is a Riemannian manifold with two Riemannian metric tensors $g_{a}$ and $g_{b}$ such that
(2.9) $g_{a}(V,V)\leq g_{b}(V,V)\quad\forall V\in TM$
then the corresponding induced Riemannian distances $d_{a}=d_{g_{a}}$ and $d_{b}=d_{g_{b}}$ satisfy
(2.10) $d_{a}(x,y)\leq d_{b}(x,y)\qquad\forall x,y\in X.$
The proof of Lemma 2.3 is an easy exercise applying 2.6 and 2.7.
Remark 2.4.
As a consequence of Lemma 2.3, a Riemannian manifold with a monotone increasing sequence of metric tensors as in (1.1), has a monotone increasing sequence of distance functions as in Definition 2.2. An easy way to construct such a sequence on the sphere, ${\mathbb{S}}^{m}$, is to take a sequence of graphs over spheres, $g_{j}=g_{{\mathbb{S}}^{m}}+dh_{j}^{2}$ where $g_{{\mathbb{S}}^{m}}$ is the standard round metric on the sphere and $h_{j}:{\mathbb{S}}^{m}\to[0,1]$ has $|\nabla h_{j}|\leq|\nabla h_{j+1}|\textrm{ on }{\mathbb{S}}^{m},$ so that $L_{g_{j}}(C)\leq L_{g_{j+1}}(C)$ for all curves $C$.
2.2. Pointwise convergence
Lemma 2.5.
Suppose that $d_{j}:X\times X\to[0,\infty)$ is a monotone increasing sequence of distance functions on $X$ as in Definition 2.2 and that there is a uniform upper bound on diameter,
(2.11) $\operatorname{diam}_{j}(X)=\sup_{x,y\in X}d_{j}(x,y)\leq D_{0}\in(0,\infty)$
then there is a limit distance function:
(2.12) $d_{\infty}(x,y)=\lim_{j\to\infty}d_{j}(x,y)=\sup_{j\in{\mathbb{N}}}d_{j}(x,y)\in[0,D_{0}]$
and $(X,d_{\infty})$ is a metric space with diameter $\operatorname{diam}_{\infty}(X)\leq D_{0}$.
Proof.
We know the pointwise limit function $d_{\infty}:X\times X\to[0,D_{0}]$ exists and satisfies (2.12) by the monotone convergence theorem for sequences of real numbers. We can confirm that $d_{\infty}$ is positive definite as follows:
(2.13) $d_{\infty}(x,x)=\lim_{j\to\infty}d_{j}(x,x)=0$
and
(2.14) $d_{\infty}(x,y)=0\implies d_{j}(x,y)\leq 0\implies x=y.$
It is symmetric for all $x,y\in X$ by
(2.15) $d_{\infty}(x,y)=\lim_{j\to\infty}d_{j}(x,y)=\lim_{j\to\infty}d_{j}(y,x)=d_{\infty}(y,x)\quad\forall x,y,\in X.$
It satisfies the triangle inequality for any $x,y,z\in X$ by
(2.16) $d_{\infty}(x,y)$ $=$ $\lim_{j\to\infty}d_{j}(x,y)$
(2.17) $\leq$ $\lim_{j\to\infty}(d_{j}(x,z)+d_{j}(z,y))$
(2.18) $=$ $\lim_{j\to\infty}d_{j}(x,z)+\lim_{j\to\infty}d_{j}(z,y)$
(2.19) $=$ $d_{\infty}(x,z)+d_{\infty}(z,y)\quad\forall x,y,z\in X.$
2.3. Example without Uniform Convergence
Even if $(X,d_{j})$ is a monotone increasing sequence of compact metric spaces with a uniform upper bound on diameter pointwise convergence does not imply uniform convergence. See Example 2.6 depicted in Figure 2.1.
Figure 2.1. This figure depicts Example 2.6.
Example 2.6.
Here we adapt an example from Sormani-Wenger [SW11] as in Figure 2.1. Consider a sequence of distinct points $p_{j}\to p_{\infty}$ lying on the equator in the standard round sphere $({\mathbb{S}}^{2},g_{\mathbb{S}^{2}})$. Let $r_{j}>0$ be decreasing radii such that the balls $B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$ are disjoint and do not contain $p_{\infty}$. We define an increasing sequence of metric tensors
(2.20) $g_{j}=g_{{\mathbb{S}}^{2}}+dh_{j}^{2}$
as in Remark 2.4 inductively starting with constant $h_{0}=0$ and taking
(2.21) $h_{j+1}(x)$ $=$ $h_{j}(x)\qquad\textrm{ for }x\in{\mathbb{S}}^{2}\setminus B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$
(2.22) $=$ $H(d_{\mathbb{S}^{2}}(p_{j},x)/r_{j})\quad\textrm{ for }x\in B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$
where $H:(-1,1)\to[0,H_{0}]$ is a fixed smooth even function of compact support such that $H(0)=H_{0}>0$ and $H^{\prime}(r)\leq 0$ for $r\geq 0$.
So we have added increasingly thin wells in place of each of the balls. For all $i\leq j$, the $g_{j}$ radial distance from $p_{i}$ to any $q_{i}\in\partial B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$,
[…]
Since $a\leq\sqrt{1+a^{2}}\leq 1+a$ we have
(2.24) $H_{0}\leq d_{g_{j}}(p_{i},q)\leq r_{i}+H_{0}\quad\forall i\leq j.$
Since any point in $({\mathbb{S}}^{2},g_{j})$ outside of the wells can be joined to a pole travelling a distance $\leq\pi/2$, and poles are a distance $\pi$ apart, we can apply the triangle inequality to see that
(2.25) $\operatorname{diam}_{g_{j}}\leq(H_{0}+r_{1}+\pi/2)+(H_{0}+r_{1}+\pi/2)+\pi.$
This uniform upper bound and Lemma 2.5 implies there is a pointwise limit, $d_{j}\to d_{\infty}$. Taking the limit of (2.24) we have,
(2.26) $H_{0}\leq d_{\infty}(p_{i},p_{\infty})\quad\forall i\in{\mathbb{N}}.$
However, for $i>j$ sufficiently large
(2.27) $d_{g_{j}}(p_{i},p_{\infty})\leq d_{\mathbb{S}^{2}}(p_{i},p_{\infty})<H_{0}/2$
so there is no uniform convergence. In fact, the limit distance function $d_{\infty}$ has a sequence of disjoint $g_{j}$ balls centered at the $p_{j}$ of radius $H_{0}$, so $({\mathbb{S}}^{2},d_{\infty})$ is not compact.
3. GH Convergence to Compact Limits
Recall the following definition of the Gromov-Hausdorff distance first introduced by Edwards in [Edw75] and then rediscovered and studied extensively by Gromov in [Gro07].
Definition 3.1.
Given a pair of metric spaces, $(X_{a},d_{a})$ and $(X_{b},d_{b})$, the Gromov-Hausdorff distance between them is
(3.1) $d_{GH}\left((X_{a},d_{a}),(X_{b},d_{b})\right)=\inf\{d_{H}^{Z}\left(f_{a}(X_{a}),f_{b}(X_{b})\right)\}$
where the infimum is taken over all common metric spaces, $(Z,d_{Z})$, and over all distance preserving maps,
(3.2) $f_{a}:(X_{a},d_{a})\to(Z,d_{Z})\textrm{ and }f_{b}:(X_{b},d_{b})\to(Z,d_{Z})$
Here $d_{H}^{Z}$ denotes the Hausdorff distance,
(3.3) $d_{H}^{Z}\left(A,B\right)=\inf\{R\,:\,A\subset T_{R}(B)\textrm{ and }B\subset T_{R}(A)\}$
between subsets $A,B\subset Z$ which is defined using tubular neighborhoods of a given radius $R$,
(3.4) $T_{R}(A)=\{z\in Z\,:\,\exists p\in A\,s.t.\,d_{Z}(p,z)<R\}.$
In this section we prove the following simple theorem:
Theorem 3.2.
If $(X,d_{j})$ is a monotone increasing sequence as in Definition 2.2 with a uniform upper bound $D_{0}$ on diameter, and if $d_{j}$ converges pointwise to $d_{\infty}$ as in Lemma 2.5, and if $(X,d_{\infty})$ is a compact metric space, then $d_{j}$ converge uniformly to $d_{\infty}$ and we have Gromov-Hausdorff convergence as well.
Although this GH Convergence theorem can be proven quite easily, we will prove it through a sequence of lemmas and propositions that we will apply again to prove SWIF convergence later.
3.1. Distance on the Product Space
We introduce the following standard definition:
Definition 3.3.
Given a pair of metric spaces, $(X_{a},d_{a})$ and $(X_{b},d_{b})$, we define the taxi product metric space $(X_{a}\times X_{b},d_{sum,a,b})$ where
(3.5) $d_{sum,a,b}((x_{1},y_{1}),(x_{2},y_{2}))=d_{a}(x_{1},x_{2})+d_{b}(y_{1},y_{2})$
for any $x_{1},x_{2}\in X_{a}$ and any $y_{1},y_{2}\in X_{b}$.
Lemma 3.4.
If $(X,d_{a})$ and $(X,d_{b})$ are compact then $(X\times X,d_{sum,a,b})$ is also compact.
The proof is an exercise for the reader.
Lemma 3.5.
Given two metric spaces $(X,d_{a})$ and $(X,d_{b})$, if
(3.6) $d_{a}(x,y)\leq d_{b}(x,y)\quad\forall x,y\in X$
then the function $d_{a}:X\times X\to[0,\infty)$ is $1$-Lipschitz with respect to the distance $d_{sum,b,b}$ on $X\times X$ given in Definition 3.3. That is,
(3.7) $|d_{a}(x_{1},y_{1})-d_{a}(x_{2},y_{2})|\leq d_{sum,b,b}((x_{1},y_{1}),(x_{2},y_{2}))\qquad\forall(x_{i},y_{i})\in X\times X.$
Proof.
By the triangle inequality and then by (3.6) we have
(3.8) $|d_{a}(x_{1},y_{1})-d_{a}(x_{2},y_{2})|$$\leq$ $d_{a}(x_{1},x_{2})+d_{a}(y_{1},y_{2})$
(3.9) $\leq$ $d_{b}(x_{1},x_{2})+d_{b}(y_{1},y_{2})$
which gives our claim by (3.5) ∎
3.2. Uniform Convergence
Proposition 3.6.
Under the hypotheses of Theorem 3.2, the distance functions, $d_{j}$, converge uniformly to $d_{\infty}$ on $X\times X$.
Proof.
By Lemma 2.5, we have $d_{j}\leq d_{\infty}$ as functions on $X\times X$. By Lemma 3.5, the maps $d_{j}:X\times X\to[0,D_{0}]$ are bounded $1$-Lipschitz functions on the metric space $(X\times X,d_{sum,\infty,\infty})$, which is compact by Lemma 3.4.
By the Arzela-Ascoli Theorem, there is a subsequence, $d_{j_{k}}:X\times X\to[0,D_{0}]$, which converges uniformly to a limit function. However, we already have a pointwise limit, so that limit function equals $d_{\infty}:X\times X\to[0,D_{0}]$. That is, $\forall\epsilon>0\,\,\exists N_{\epsilon}\in\mathbb{N}\,\,s.t.\,\,\forall k\geq N_{\epsilon}$
(3.10) $d_{\infty}(x,y)-\epsilon<d_{j_{k}}(x,y)<d_{\infty}(x,y)+\epsilon\qquad\forall x,y\in X\times X.$
By the assumption of monotonicity, for all $j\geq j_{k}\geq j_{N_{\epsilon}}$,
(3.11) $d_{\infty}(x,y)-\epsilon<d_{j_{k}}(x,y)\leq d_{j}(x,y)\leq d_{\infty}(x,y)\qquad\forall x,y\in X\times X.$
3.3. Constructing a Common Metric Space
Recall that the definition of the Gromov-Hausdorff distance involves a common metric space $Z$ and distance preserving maps [Definition 3.1]. The following lemma is not part of the original proof by Gromov that uniform convergence implies Gromov-Hausdorff convergence. Here we create a different common metric space $Z$ that allows us to prove intrinsic flat convergence later as well. It is a simplification of the $Z$ constructed by Allen-Perales-Sormani in [APS24].
Proposition 3.7.
Given two metric spaces $(X,d_{a})$ and $(X,d_{b})$ and an $\epsilon>0$ such that
(3.12) $d_{b}(x,y)-\epsilon\leq d_{a}(x,y)\leq d_{b}(x,y)\quad\forall x,y\in X$
then there exists a metric space
(3.13) $(Z_{a,b}=X\times[0,h],d_{Z_{a,b}})\textrm{ where }h=\epsilon/2,$
with the distance between $z_{1}=(x_{1},t_{1})$ and $z_{2}=(x_{2},t_{2})$ defined by
(3.14) $d_{Z_{a,b}}(z_{1},z_{2})=\min\{d_{taxi_{b}}(z_{1},z_{2}),(h-t_{1})+(h-t_{2})+d_{a}(x_{1},x_{2})\}.$
[…]
so that the identity map,
(3.16) $id_{a,b}:(Z_{a,b},d_{taxi_{b}})\to(Z_{a,b},d_{Z_{a,b}})$
is $1$-Lipschitz and there are distance preserving maps:
[…]
so that
(3.19) $d_{GH}((X,d_{a}),(X,d_{b}))\leq d_{H}^{Z}(f_{a}(X),f_{b}(X))\leq h.$
The intuition behind this construction is that we are taking the taxi product of $(X,d_{b})$ with the interval $([0,h],d_{\mathbb{R}})$ and then gluing $(X,d_{a})$ to this taxi product at $t=h$ which gives possibly shorter distances. See Figure 3.1.
Figure 3.1. On the left we see $(X_{a},d_{a})$ above $(X_{b},d_{b})$ mapping into $(Z_{a,b},d_{Z_{a,b}})$ and on the right we see two paths between points in $Z_{a,b}$ where the first achieves $d_{taxi_{b}}$ and the second takes a short cut through the image of $X_{a}$. The shorter of the two paths achieves the minimum in the definition of $d_{Z_{a,b}}$.
Proof.
First we confirm that $d_{Z_{a,b}}$ is a metric. It is clearly definite and symmetric. To see that it satisfies the triangle inequality we have several cases. Let $z_{i}=(x_{i},t_{i})\in Z_{a,b}$, $i=1,2,3$ and
(3.20) $D_{13}=d_{Z_{a,b}}(z_{1},z_{2})+d_{Z_{a,b}}(z_{2},z_{3}).$
Case I: Assume that $d_{Z_{a,b}}(z_{1},z_{2})=d_{taxi_{b}}(z_{1},z_{2})$ and $d_{Z_{a,b}}(z_{2},z_{3})=d_{taxi_{b}}(z_{2},z_{3})$ Then since $d_{taxi_{b}}$ is a distance function, and thus satisfies the triangle inequality, and by definition of $d_{Z_{a,b}}$, we get
(3.21) $D_{13}=d_{taxi_{b}}(z_{1},z_{2})+d_{taxi_{b}}(z_{2},z_{3})\geq d_{taxi_{b}}(z_{1},z_{3})\geq d_{Z_{a,b}}(z_{1},z_{3}).$
Case II: Assume that we have $d_{Z_{a,b}}(z_{1},z_{2})=d_{taxi_{b}}(z_{1},z_{2})$ and
(3.22) $d_{Z_{a,b}}(z_{2},z_{3})=(h-t_{2})+(h-t_{3})+d_{a}(x_{2},x_{3}).$
Then by $d_{b}\geq d_{a}$, triangle inequality in $\mathbb{R}$ and for $d_{a}$, we get
$D_{13}$ $=d_{taxi_{b}}(z_{1},z_{2})+(h-t_{2})+(h-t_{3})+d_{a}(x_{2},x_{3})$
$\geq d_{a}(x_{1},x_{2})+|t_{1}-t_{2}|+(h-t_{2})+(h-t_{3})+d_{a}(x_{2},x_{3})$
$\geq d_{a}(x_{1},x_{2})+(h-t_{1})+(h-t_{3})+d_{a}(x_{2},x_{3})$
$\geq d_{a}(x_{1},x_{3})+(h-t_{1})+(h-t_{3})\geq d_{Z_{a,b}}(z_{1},z_{3}).$
Case III: Assume that
(3.23) $d_{Z_{a,b}}(z_{1},z_{2})=(h-t_{1})+(h-t_{2})+d_{a}(x_{1},x_{2})$
(3.24) $d_{Z_{a,b}}(z_{2},z_{3})=(h-t_{2})+(h-t_{3})+d_{a}(x_{2},x_{3}).$
Then, since $h-t_{2}\geq 0$ and triangle inequality for $d_{a}$, we get
(3.25) $D_{13}$ $=(h-t_{1})+(h-t_{2})+d_{a}(x_{1},x_{2})$
(3.26) $\qquad+(h-t_{2})+(h-t_{3})+d_{a}(x_{2},x_{3})$
(3.27) $\geq(h-t_{1})+d_{a}(x_{1},x_{2})+(h-t_{3})+d_{a}(x_{2},x_{3})$
(3.28) $\geq d_{a}(x_{1},x_{3})+(h-t_{1})+(h-t_{3})$
(3.29) $\geq d_{Z_{a,b}}(z_{1},z_{3}),$
which completes the proof of the triangle inequality on $(Z_{a,b},d_{Z_{a,b}})$.
We have the claimed $1$-Lipschitz identity map, because $d_{Z_{a,b}}\leq d_{taxi_{b}}$. We have the claims that $f_{a}$ and $f_{b}$ are distance preserving maps as follows. By the definition of $d_{Z_{a,b}}$, $f_{a}$ and $d_{b}\geq d_{a}$,
$d_{Z_{a,b}}(f_{a}(x),f_{a}(y))=d_{Z_{a,b}}((x,h),(y,h))=\min\{d_{b}(x,y),d_{a}(x,y)\}=d_{a}(x,y).$
By definition of $d_{Z_{a,b}}$, $f_{b}$, (3.12) and $h=\epsilon/2$ we have
(3.30) $d_{Z_{a,b}}(f_{b}(x),f_{b}(y))=\min\{d_{b}(x,y),2h+d_{a}(x,y)\}=d_{b}(x,y).$
Furthermore, by the definition of $d_{Z_{a,b}}$, for any $x\in X$, it holds
(3.31) $d_{Z_{a,b}}((x,0),(x,h))=h,$
which implies that for any $R>h$,
(3.32) $f_{b}(X)\subset T_{R}(f_{a}(X))\textrm{ and }f_{a}(X)\subset T_{R}(f_{b}(X)),$
so the Hausdorff distance $d_{H}^{Z}(f_{a}(X),f_{b}(X))\leq h$ which gives our final claim. ∎
3.4. Proof of GH Convergence
Theorem 3.2 is now easy to prove:
Proof.
First we apply Prop 3.6 to get uniform convergence from below. Then we apply Proposition 3.7 to $(X,d_{j})$ and $(X,d_{\infty})$ to see that
(3.33) $d_{GH}((X,d_{j}),(X,d_{\infty}))<\epsilon_{j}\to 0.$
Remark 3.8.
Note that in the proof of Theorem 3.2 it was essential that we assumed $(X,d_{\infty})$ is compact, as seen in Example 2.6. This example has the monotonicity but no Gromov-Hausdorff limit. In fact Gromov proved that if a sequence of compact metric spaces has a GH limit then the entire sequence can be embedded via distance preserving maps into a common metric space, $(Z,d_{Z})$ [Gro81]. However, the Riemannian manifolds in Example 2.6 have an unbounded number of disjoint balls of radius $1$, so this is impossible.
3.5. A Common Compact Metric Space for the Whole Sequence
Gromov proved in [Gro81] that an entire GH-converging sequence of metric spaces can be embedded with distance preserving maps into a common compact metric space $Z$ along with the limit so that the sequence of images converges in the Hausdorff sense inside $Z$ to the image of the limit. Here we construct a specific $Z$ which we will use later to prove SWIF convergence and Theorem 1.1. See Figure 3.2.
Proposition 3.9.
If $(X,d_{j})$ is uniformly converging to compact $(X,d_{\infty})$
(3.34) $d_{\infty}(x,y)-\epsilon_{j}<d_{j}(x,y)\leq d_{\infty}(x,y)\quad\forall x,y\in X$
with $\epsilon_{j}$ decreasing to 00 then taking $Z_{j}$ of Proposition 3.7, we define
(3.35) $Z=\bigsqcup_{j=1}^{\infty}Z_{j}\,\,|_{\sim}\textrm{ where }Z_{j}=Z_{j,\infty}=X_{j}\times[0,h_{j}]$
where $(x,0)\in Z_{j}$ is identified with $(x,0)\in Z_{k}$ for all $x\in X$ and all $j,k\in{\mathbb{N}}$. Then $d_{Z}:Z\times Z\to[0,\infty)$ given by
(3.36) $d_{Z}(z_{j},z_{k})=\inf\{d_{Z_{j}}((x_{j},t_{j}),(x,0))+d_{Z_{k}}((x,0),(x_{k},t_{k}))\,|\,x\in X\}$
for
(3.37) $z_{j}=(x_{j},t_{j})\in Z_{j}\textrm{ and }z_{k}=(x_{k},t_{k})\in Z_{k}\qquad j\neq k,$
and, otherwise,
(3.38) $d_{Z}(z_{j},z_{k})=d_{Z_{j}}(z_{j},z_{k}),$
so the inclusion maps
(3.39) $\zeta_{j}:(Z_{j},d_{Z_{j}})\to(Z,d_{Z})$
are distance preserving is a distance function. In addition, taking
(3.40) $f_{j}=f_{j,(j,\infty)}:(X,d_{j})\to Z_{j}\textrm{ and }f_{\infty,(j,\infty)}:(X,d_{\infty})\to Z_{j}$
from Proposition 3.7, we have distance preserving maps:
(3.41) $\chi_{j}=\zeta_{j}\circ f_{j}:(X,d_{j})\to(Z,d_{Z})$
and
(3.42) $\chi_{\infty}=\zeta_{j}\circ f_{\infty,(j,\infty)}:(X,d_{\infty})\to(Z,d_{Z}),$
where $\chi_{\infty}(x)=\zeta_{j}(x,0)$ does not depend on $j$ and is an isometry onto,
(3.43) $Z_{0}=\chi_{\infty}(X)=\zeta_{j}(f_{\infty,(j,\infty)}(X))\subset Z.$
Furthermore, we have
(3.44) $d_{Z}(\chi_{j}(x),\chi_{\infty}(x))=d_{Z_{j}}((x,h_{j}),(x,0))=h_{j}$
and
(3.45) $d_{H}^{Z}(\chi_{j}(X)),\chi_{\infty}(X))\leq h_{j}.$
Finally $(Z,d_{Z})$ is compact.
Intuitively we have glued the $Z_{j}$ together like the pages of a book, $\zeta_{j}:Z_{j}\to Z$, along the spine, $Z_{0}$, as in Figure 3.2.
Figure 3.2. On the left, we see the sequence $(X,d_{j})$ above with downward maps $f_{j,(j,\infty)}:(X,d_{j})\to(Z_{j,\infty},d_{Z_{j,\infty}})$ and upward maps $f_{\infty,(j,\infty)}:(X,d_{\infty})\to(Z_{j,\infty},d_{Z_{j,\infty}})$ as constructed in Proposition 3.7 and depicted in Figure 3.1. On the right, we glue together all the $(Z_{j,\infty},d_{Z_{j,\infty}})$ to create $(Z,d_{Z})$ as in Proposition 3.9 with maps $\chi_{j}:X_{j}\to Z$ and $\chi_{\infty}:X_{j}\to Z$.
Proof.
It is clear that $d_{Z}$ is symmetric and positive definite. To see that it satisfies the triangle inequality we have several cases. Let $z_{j_{i}}=(x_{i},t_{i})\in Z_{j_{i}}$, $j_{i}\in\mathbb{N}$ and $i=1,2,3$,
(3.46) $D_{13}=d_{Z}(z_{j_{1}},z_{j_{2}})+d_{Z}(z_{j_{2}},z_{j_{3}}).$
Case I: Assume that $j_{1}=j_{2}$ and $j_{2}=j_{3}$. In this case, $d_{Z}=d_{Z_{j_{i}}}$ which satisfies the triangle inequality since $d_{Z_{j_{i}}}$ is a distance function.
Case II: Assume that $j_{1}=j_{2}$ and $j_{2}\neq j_{3}$. By $j_{1}=j_{2}$, the triangle inequality for $d_{Z_{j_{2}}}$ and definition of $d_{Z}$ we get
(3.47) $D_{13}=$ $d_{Z_{j_{2}}}(z_{j_{1}},z_{j_{2}})+\inf_{x\in X}\{d_{Z_{j_{2}}}(z_{j_{2}},(x,0))+d_{Z_{j_{3}}}((x,0),z_{j_{3}})\}$
(3.48) $=$ $\inf_{x\in X}\{d_{Z_{j_{2}}}(z_{j_{1}},z_{j_{2}})+d_{Z_{j_{2}}}(z_{j_{2}},(x,0))+d_{Z_{j_{3}}}((x,0),z_{j_{3}})\}$
(3.49) $\geq$ $\inf_{x\in X}\{d_{Z_{j_{2}}}(z_{j_{1}},(x,0))+d_{Z_{j_{3}}}((x,0),z_{j_{3}})\}=d_{Z}(z_{j_{1}},z_{j_{3}}).$
Case III: Assume that $j_{1}\neq j_{2}$ and $j_{2}\neq j_{3}$. Take $x,x^{\prime}\in X$, using the triangle inequality for $d_{Z_{j_{2}}}$ and $d_{Z_{j_{1}}}$, and the definition of $d_{Z}$ we get
$d_{Z_{j_{1}}}(z_{j_{1}},(x,0))+d_{Z_{j_{2}}}((x,0),z_{j_{2}})+d_{Z_{j_{2}}}(z_{j_{2}},(x^{\prime},0))+d_{Z_{j_{3}}}((x^{\prime},0),z_{j_{3}})$
$\geq d_{Z_{j_{1}}}(z_{j_{1}},(x,0))+d_{Z_{j_{2}}}((x,0),(x^{\prime},0))+d_{Z_{j_{3}}}((x^{\prime},0),z_{j_{3}})$
$\geq d_{Z_{j_{1}}}(z_{j_{1}},(x,0))+d_{Z_{j_{1}}}((x,0),(x^{\prime},0))+d_{Z_{j_{3}}}((x^{\prime},0),z_{j_{3}})$
$\geq d_{Z_{j_{1}}}(z_{j_{1}},(x^{\prime},0))+d_{Z_{j_{3}}}((x^{\prime},0),z_{j_{3}})\geq d_{Z}(z_{j_{1}},z_{j_{3}}).$
Taking the infimum over $x,x^{\prime}\in X$ in the previous expression we obtain $D_{13}\geq d_{Z}(z_{j_{1}},z_{j_{3}})$. This concludes the proof that $(Z,d_{Z})$ is a metric space.
The inclusion maps $\zeta_{j}$ are distance preserving by their definition and the definition of $d_{Z}$. Since by Proposition 3.7 $f_{j}=f_{j,(j,\infty)}$ and $f_{\infty,(j,\infty)}$ are distance preserving maps, then the compositions $\chi_{j}=\zeta_{j}\circ f_{j}$ and $\chi_{\infty}=\zeta_{j}\circ f_{\infty,(j,\infty)}$ are distance preserving maps. Since we are identifying points of the form $(x,0)\in Z_{j}$ with points $(x,0)\in Z_{k}$ for all $x\in X$, then $\chi_{\infty}(x)=\zeta_{j}(x,0)$ does not depend on $j$. Furthermore, $\chi_{\infty}$ is an isometry onto $Z_{0}=\chi_{\infty}(X)$.
Since $\chi_{j}$ and $\chi_{\infty}$ are distance preserving maps and by (3.19) in Proposition 3.7 we have
(3.50) $d_{Z}(\chi_{j}(x),\chi_{\infty}(x))\leq d_{Z_{j}}((x,h_{j}),(x,0))\leq h_{j}$
so we have (3.44)-(3.44).
Finally we check that $(Z,d_{Z})$ is compact. Given any sequence $z_{i}\in Z$,
(3.51) $\exists\,j_{i}\in{\mathbb{N}}\,\,\exists\,(x_{i},t_{i})\in Z_{j_{i}}=X\times[0,h_{j_{i}}]\textrm{ such that }z_{i}=\zeta_{j_{i}}(x_{i},t_{i}).$
Since $(X,d_{\infty})$ is compact, a subsequence $x_{i}\to x_{\infty}$. Since $[0,1]$ is compact, a further subsequence can be taken such that $t_{i}\to t_{\infty}$. A further subsequence can be taken to guarantee that either $j_{i}\to\infty$ or $j_{i}$ is constant $j_{i}=j_{0}$. In the diverging case we use $t_{i}\leq h_{j_{i}}\to 0$ to see that $t_{\infty}=0$, thus our subsequence converges to
(3.52) $z_{\infty}=\chi_{\infty}(x_{\infty})=\zeta_{j_{i}}(x_{\infty},0)\in Z_{0}$
which we see as follows:
(3.53) $d_{Z}(z_{i},z_{\infty})$ $=$ $d_{Z}(\zeta_{j_{i}}(x_{i},t_{i}),\zeta_{j_{i}}(x_{\infty},0))$
(3.54) $=$ $d_{Z_{j_{i}}}((x_{i},t_{i}),(x_{\infty},0))$
(3.55) $\leq$ $d_{\infty}(x_{i},x_{\infty})+|t_{i}-0|\to 0.$
In the constant, $j_{i}=j_{0}$ case, our subsequence converges to
(3.56) $z_{\infty}=\zeta_{j_{0}}(x_{\infty},t_{\infty})$
which we see as follows:
(3.57) $d_{Z}(z_{i},z_{\infty})$ $=$ $d_{Z}(\zeta_{j_{0}}(x_{i},t_{i}),\zeta_{j_{0}}(x_{\infty},t_{\infty}))$
(3.58) $=$ $d_{Z_{j_{0}}}((x_{i},t_{i}),(x_{\infty},t_{\infty}))$
(3.59) $\leq$ $d_{\infty}(x_{i},x_{\infty})+|t_{i}-t_{\infty}|\to 0.$
Thus $Z$ is compact. ∎
Lemma 3.10.
Assume the hypotheses of Proposition 3.9. For all $j\in{\mathbb{N}}$ there exists
(3.60) $W_{j}=\bigcup_{k=j}^{\infty}\chi_{k}(Z_{k})\subset Z$
and there is a $1$-Lipschitz map, $F_{j}:(W_{j},d_{Z})\to(X,d_{j})$ where
(3.61) $F_{j}(\zeta_{k}(x,t))=x\quad\forall\,k\geq j\,\forall x\in X,\,\forall t\in[0,h_{k}].$
Proof.
Let $z_{1},z_{2}\in W_{j}\subset Z$. Then
(3.62) $\exists k_{i}\geq j,\,\,\exists x_{i}\in X\,\,\exists t_{i}\in[0,h_{k_{i}}]\textrm{ s.t. }z_{i}=\chi_{k_{i}}(x_{i},t_{i}).$
If $k_{1}=k_{2}$, since $d_{\infty}\geq d_{k_{i}}\geq d_{j}$, by the minimum in the definition of $d_{Z_{k_{i}}}$ and the monotonicity of the distance functions, we get
(3.63) $d_{Z}(z_{1},z_{2})\geq d_{k_{i}}(x_{1},x_{2})\geq d_{j}(x_{1},x_{2}).$
If $k_{2}\neq k_{1}$, then observe that for all $x\in X$,
(3.64) $d_{Z_{k_{i}}}(z_{i},(x,0))=d_{Z_{k_{i}}}((x_{i},t_{i}),(x,0))\geq d_{k_{i}}(x_{i},x)\geq d_{j}(x_{i},x)$
by the definition of $d_{Z_{k_{i}}}$ because $t_{i}\geq 0$ and $h_{k_{i}}-t_{i}+h_{k_{i}}\geq 0$, followed by the monotonicity of the distance functions. Thus
(3.65) $d_{Z_{k_{1}}}(z_{1},(x,0))+d_{Z_{k_{2}}}((x,0),z_{2})\geq d_{j}(x_{1},x)+d_{j}(x_{2},x).$
Taking the infimum over all $x\in X$ on the left and applying the triangle inequality for $d_{j}$ on the right we have,
(3.66) $d_{Z}(z_{1},z_{2})\geq d_{j}(x_{1},x_{2}).$
4. Intrinsic Flat Convergence
In this section we review the definition of integral current spaces which include all oriented Riemannian manifolds of finite volume with boundary of finite volume. We then review the Sormani-Wenger Intrinsic Flat (SWIF) distance between these spaces. Finally we state and prove Theorem 4.16 concerning the convergence of monotone increasing sequences of integral current spaces which implies the SWIF convergence claimed in Theorem 1.1.
4.1. Lipschitz Functions and Tuples on Metric Spaces
We say that a function, $F:(X,d_{X})\to(Y,d_{Y})$ is $K$-Lipschitz if
(4.1) $d_{Y}(F(x_{1},x_{2}))\leq Kd_{X}(x_{1},x_{2})\quad\forall x_{1},x_{2}\in X$
We define the Lipschitz constant of $F$ to be the smallest such $K$.
Lemma 4.1.
Suppose that $(X,d_{a})$ and $(X,d_{b})$ are two metric spaces and
(4.2) $d_{a}(x_{1},x_{2})\leq d_{b}(x_{1},x_{2})\quad\forall x_{1},x_{2}\in X.$
If a function $\pi:X\to Y$ is $K$-Lipschitz as a map $\pi:(X,d_{a})\to(Y,d_{Y})$ then it is also $K$-Lipschitz as a map $\pi:(X,d_{b})\to(Y,d_{Y})$.
Proof.
(4.3) $d_{Y}(\pi(x_{1}),\pi(x_{2}))\leq Kd_{a}(x_{1},x_{2})\leq Kd_{b}(x_{1},x_{2})\quad\forall x_{1},x_{2}\in X.$
By the definition in [AK00] an m-tuple, $(\pi_{0},\pi_{1},...,\pi_{m})$, of Lipschitz functions with respect to $d_{a}$ consists of a bounded Lipschitz function, $\pi_{0}:(X,d_{a})\to{\mathbb{R}}$, and $m$ Lipschitz functions, $\pi_{k}:(X,d_{a})\to{\mathbb{R}}$, for $k=1,\dots,m$.
Corollary 4.2.
Suppose that $(X,d_{a})$ and $(X,d_{b})$ are two metric spaces and
(4.4) $d_{a}(x,y)\leq d_{b}(x,y)\quad\forall x,y\in X.$
Then tuples of Lipschitz functions with respect to $d_{a}$ are tuples of Lipschitz functions with respect to $d_{b}$.
Lemma 4.3.
Given the hypotheses of Proposition 3.9 and Lemma 3.10. If $\pi:(X,d_{j})\to Y$ is Lipschitz $K_{j}$, and $F_{j}:(W_{j},d_{Z})\to(X,d_{j})$ of Lemma 3.10, then
(4.5) $\tilde{\pi}:(W_{j},d_{Z})\to Y\textrm{ such that }\tilde{\pi}(z)=\pi\circ F_{j}(z)$
is also Lipschitz $K_{j}$ and satisfies
(4.6) $\tilde{\pi}(\chi_{\infty}(x))=\pi(x)\quad\,\forall x\in X.$
Remark 4.4.
Note that if we try to define $\tilde{\pi}$ as a function from all of $(Z,d_{Z})$ to ${\mathbb{R}}$ as in (4.5) then it might not be Lipschitz. The proof depends strongly on the restriction to $W_{j}\subset Z$ and Lemma 3.10.
Proof.
Recall $F_{j}:(W_{j},d_{Z})\to(X,d_{j})$ are Lipschitz one functions, so
(4.7) $d_{Y}(\tilde{\pi}(z_{1}),\tilde{\pi}(z_{2}))\leq K_{j}\,d_{j}(F_{j}(z_{1}),F_{j}(z_{2}))\leq K_{j}\,d_{Z}(z_{1},z_{2})\quad\forall z_{1},z_{2}\in Z.$
4.2. Charts into our Metric Space
This is needed to understand integral structures on $(X,d_{j})$ vs $(X,d_{j+1})$ and $(X,d_{\infty})$.
Lemma 4.5.
Suppose $(X,d_{a})$ and $(X,d_{b})$ are two metric spaces and
(4.8) $d_{a}(x,y)\leq d_{b}(x,y)\quad\forall x,y\in X.$
If $\psi:U\subset{\mathbb{R}}^{m}\to X$ is a Lipschitz chart with respect to $d_{b}$ on $X$, then it is Lipschitz with respect to $d_{a}$ on $X$.
Proof.
We have a Lipschitz constant $K_{\psi}\in(0,\infty)$ such that
(4.9) $d_{b}(\psi(a_{1}),\psi(a_{2}))\leq K_{\psi}|a_{1}-a_{2}|\qquad\forall a_{1},a_{2}\in U.$
By (4.8), we have
(4.10) $d_{a}(\psi(a_{1}),\psi(a_{2}))\leq K_{\psi}|a_{1}-a_{2}|\qquad\forall a_{1},a_{2}\in U.$
Remark 4.6.
Sometimes in the definition of a rectifiable space or current it is written that the charts $\psi_{i}:U_{i}\to X$ are bi-Lipschitz. This is true by Lemma 4 in Kirchheim’s [Kir94] as soon as one has a countable collection of Lipschitz charts (cf. Lemma 4.1 of [AK00]). The bi-Lipschitz charts are not necessarily the same collection of charts. So even if the charts $\psi_{i}:(U_{i}\subset{\mathbb{R}}^{m},d_{\mathbb{R}^{m}})\to(X,d_{b})$ are bi-Lipschitz onto their images, they are only Lipschitz as charts into $(X,d_{a})$ not bi-Lipschitz, and so one would need to apply Kirchheim’s theorem to produce bi-Lipschitz charts into $(X,d_{b})$.
4.3. Ambrosio-Kirchheim’s integral currents on complete metric spaces
Here we review the work of Ambrosio-Kirchheim in [AK00]. They defined $m$ dimensional currents on complete metric spaces, $(Z,d_{Z})$ as multilinear functionals, $T$, acting on Lipschitz tuples satisfying various hypotheses [AK00]. For example, given a compact $m$-dimensional oriented Riemannian manifold, $(M,g)$, we can define the current,
[…]
which is well defined because Lipschitz maps are differentiable almost everywhere. The integration is actually defined by integrating over atlas of disjoint oriented charts and taking the sum.
An $m$ dimensional integer rectifiable current on a complete metric space, $(Z,d_{Z})$, is a multilinear functional, $T$, defined on Lipschitz tuples on $(Z,d_{Z})$ that has parametrization which is an atlas of Lipschitz charts
(4.12) $\psi_{i}:(U_{i}\subset{\mathbb{R}}^{m},d_{{\mathbb{R}}^{m}}\to(Z,d_{Z})$
defined on Borel sets, $U_{i}\subset{\mathbb{R}}^{m}$, with integer multiplicities $a_{i}$, such that total Hausdorff measures satisfies
(4.13) $\sum_{i=1}^{\infty}|a_{i}|\mathcal{H}_{d_{Z}}^{m}(\psi_{i}(U_{i}))<\infty.$
The integer rectifiable current $T$ is defined by the weighted sum:
(4.14) $T=\sum_{i=1}^{\infty}\,a_{i}\,\psi_{i\#}[[U_{i}]],$
where
(4.15) $\psi_{\#}[[U]](\pi_{0},\pi_{1},...,\pi_{m})=\int_{U}(\pi_{0}\circ\psi)d(\pi_{1}\circ\psi)\wedge\cdots\wedge d(\pi_{m}\circ\psi).$
The integral is well defined because charts are Lipschitz and thus differentiable almost everywhere. The weighted sum of the integrals defining $T(\pi_{0},...,\pi_{m})$ is finite because the weighted sum of the Hausdorff measures in (4.13), $\max\{\pi_{0}\}$, and the product $\operatorname{Lip}(\pi_{j})$ for $j=1,\ldots,m$ can be used to bound the integrals. Warning: the Lipschitz charts and the Hausdorff measures depend on the distance function on $Z$. See Lemma 4.7 below.
By Stoke’s Theorem, an oriented Riemannian manifold of dimension $m$, $M$, has
[…]
So it is natural to define the boundary of a current by
(4.17) $\partial T(\pi_{1},...,\pi_{m})=T(1,\pi_{1},...,\pi_{m}).$
An integral current is an integer rectifiable current whose boundary is also integer rectifiable. Be warned however that the parametrization of the boundary, $\partial T$, is not necessarily found by taking the boundaries of the atlas of charts parametrizing, $T$.
The mass measure, $||T||_{d_{Z}}$, of a current $T$ on a complete metric space, $(Z,d_{Z})$, is the smallest finite Borel measure such that
(4.18) $T(\pi_{0},\pi_{1},...,\pi_{m})\leq\prod_{i=1}^{m}\operatorname{Lip}(\pi_{i})\,\,\int_{Z}|\pi_{0}|\,\,||T||_{d_{Z}}$
for all tuples and the mass of $T$ is
(4.19) ${\mathbf{M}}_{d_{Z}}(T)=||T||_{d_{Z}}(Z).$
The support of a current is
(4.20) $\operatorname{spt}_{d_{Z}}(T)=\{z\in Z\,:\,||T||_{d_{Z}}(B_{d_{Z}}(z,r))>0\,\,\forall r>0\}$
and the set of an $m$ dimensional current is
(4.21) $\operatorname{set}_{d_{Z}}(T)=\left\{z\in Z\,:\,\liminf_{r\to 0}||T||_{d_{Z}}(B_{d_{Z}}(z,r))/r^{m}>0\,\,\forall r>0\right\}.$
Ambrosio-Kirchheim prove that $\operatorname{set}_{d_{Z}}(T)$ is rectifiable and that the closure of this set is the support of $T$,
(4.22) $\operatorname{spt}(T)=\overline{\operatorname{set}_{d_{Z}}(T)}.$
Given a Lipschitz map, $F:(X,d_{X})\to(Y,d_{Y})$, the pushforward of an integral current, $T$, on $(X,d_{X})$, to an integral current, $F_{\#}(T)$, on $(Y,d_{Y})$ is defined by
(4.23) $F_{\#}(T)(\pi_{0},...,\pi_{m})=T(\pi_{0}\circ F,...,\pi_{m}\circ F)$
for any Lipschitz tuple on $(Y,d_{Y})$. To see that $F_{\#}(T)$ is an integral current, Ambrosio-Kirchheim prove that the mass is bounded as follows:
(4.24) ${\mathbf{M}}_{d_{Y}}(F_{\#}(T))\leq(\operatorname{Lip}(F))^{m}{\mathbf{M}}_{d_{X}}(T).$
For our purposes we need the following lemma.
Lemma 4.7.
Suppose $(X,d_{a})$ and $(X,d_{b})$ are two metric spaces and
(4.25) $d_{a}(x,y)\leq d_{b}(x,y)\quad\forall x,y\in X.$
If $S$ is an integral current on $(X,d_{b})$ then it is an integral current on $(X,d_{a})$ defined using the same collection of Lipschitz charts and weights with
(4.26) ${\mathbf{M}}_{d_{a}}(S)\leq{\mathbf{M}}_{d_{b}}(S)\textrm{ and }{\mathbf{M}}_{d_{a}}(\partial S)\leq{\mathbf{M}}_{d_{b}}(\partial S).$
The mass measures also satisfy
(4.27) $||S||_{d_{a}}\leq||S||_{d_{b}}\textrm{ and }||\partial S||_{d_{a}}\leq||\partial S||_{d_{b}}.$
Proof.
By the definition of integral current structure on $(X,d_{b})$, there is a countable collection of $d_{b}$-Lipschitz charts, $\psi_{i}:U_{i}\subset{\mathbb{R}}^{m}\to X$, and integers $a_{i}$ such that
(4.28) $S(\omega)=\sum_{i=1}^{\infty}a_{i}\psi_{i\#}[[U_{i}]](\omega)$
for any $m$ tuple, $\omega$, of $d_{b}$-Lipschitz functions. By Lemma 4.5 these charts are also $d_{a}$-Lipschitz charts, so the weighted sum in (4.28) is well defined acting on any $m$ tuple, $\omega$, of $d_{a}$-Lipschitz functions. Note that the same can be done for $\partial S$. Thus $S$ is an integral current on $(X,d_{a})$.
To estimate the inequalities of the mass measures, recall that in [AK00] the mass measure of a current, $S$, on a metric space $(X,d)$ is defined as the smallest Borel measure, $\mu$, on $X$ such that
[…]
for all tuples, $(\pi_{0},\pi_{1},...,\pi_{m})$, of $d$-Lipschitz functions. By Corollary 4.2, for any tuple, $(\pi_{0},\pi_{1},...,\pi_{m})$, of $d_{a}$-Lipschitz functions, we also have (4.29) with $\mu=||S||_{d_{b}}$. Thus the smallest measure for the tuples of $d_{a}$-Lipschitz functions exists and is smaller,
(4.30) $||S||_{d_{a}}\leq||S||_{d_{b}}.$
The same can be done for $\partial S$. This also gives us the total mass estimate in (4.26). ∎
Note in particular the above lemma implies that distance preserving maps push forward currents conserving their masses, boundaries, and boundary masses.
Remark 4.8.
The converse of Lemma 4.7 does not hold. If $T$ is an integral current on $(X,d_{a})$, it does not necessarily define an integral current on $(X,d_{b})$ for $d_{b}\geq d_{a}$. It is a multilinear functional acting on tuples in $(X,d_{b})$ because these are also tuples on $(X,d_{a})$ by Lemma 4.1. However it does not necessarily have finite mass. In addition, a parametrization of $T$ by Lipschitz charts into $(X,d_{a})$ might not give Lipschitz charts into $(X,d_{b})$.
Soon we will be considering a sequence of metric spaces $(X,d_{j})$ with an integral current $T$, and the limit, $(X,d_{\infty})$. Since $d_{\infty}\geq d_{j}$, we will need to overcome the lack of a converse to Lemma 4.7. We will see later in the proof of Lemma 4.11, that we can sometimes try to rewrite a parametrization of an integral current on $(X,d_{j})$ to obtain charts that are also Lipschitz into $(X,d_{\infty})$. We will see later in the proof of Theorem 4.16, which is needed to complete the proof of Theorem 1.1, that we can sometimes find an integral current, $T_{\infty}$, on $(X,d_{\infty})$ and later show $T_{\infty}=T$ in the following sense:
Definition 4.9.
If we have an integral current $T$ on $(X,d_{a})$ and there is an integral current, $S$, on $(X,d_{b})$ where $d_{b}\geq d_{a}$ such that
(4.31) $S(\pi_{0},\pi_{1},...,\pi_{m})=T(\pi_{0},\pi_{1},...,\pi_{m}).$
for any tuple of $d_{a}$ Lipschitz functions on $X$, then we say that $S=T$ viewed as an integral current on $(M,d_{a})$. This means we can use any parametrizations of $S$ and $\partial S$ with charts into $(X,d_{b})$ as parametrizations for $T$ and $\partial T$ with the same charts and weights into $(X,d_{a})$ respectively. So we can view $T$ as an integral current on $(X,d_{b})$.
4.4. Weak Convergence of Currents
A sequence $T_{j}$ converges weakly as currents in $(Z,d_{Z})$ to an integral current, $T_{\infty}$ iff
(4.32) $T_{j}(\pi_{0},\pi_{1},...,\pi_{m})\to T_{\infty}(\pi_{0},\pi_{1},...,\pi_{m})$
for all tuples, $(\pi_{0},\pi_{1},...,\pi_{m})$ on $(Z,d_{Z})$.
We will apply the following powerful theorem of Ambrosio-Kirchheim:
Theorem 4.10 (Ambrosio-Kirchheim).
If $Z$ is compact and $T_{j}$ have uniform upper bounds
(4.33) ${\mathbf{M}}(T_{j})\leq V_{0}\textrm{ and }{\mathbf{M}}(\partial T_{j})\leq A_{0}$
then a subsequence $T_{j_{k}}$ converges weakly to an integral current, $T_{\infty}$, and the mass measure is lower semicontinuous:
(4.34) $||T_{\infty}||(U)\leq\liminf_{k\to\infty}||T_{j_{k}}||(U)$
for any open set $U\subset Z$.
4.5. Integral Current Spaces
In [SW11], Sormani and Wenger define an $m$-dimensional integral current space, $(X,d,T)$, as a metric space, $(X,d)$, with an integral current, $T$, defined on the closure of $(X,d_{X})$ such that the $\operatorname{set}(T)=X$ where
(4.35) $\operatorname{set}(T)=\left\{x\in\bar{X}\,:\,\liminf_{r\to 0}||T||_{d}(B_{d}(x,r))/r^{m}>0\right\}.$
Such metric spaces $(X,d)$ are rectifiable and parametrized by the charts that parametrize their integral current structure, $T$.
Given a smooth compact oriented $m$-dimensional Riemannian manifold with boundary, $(M,g)$, it can be viewed as an integral current space, $(M,d_{g},[[M]])$. Here the mass measure is the volume measure,
(4.36) ${\mathbf{M}}_{d_{g}}(M)=\operatorname{vol}_{g}(M),$
and
(4.37) $\operatorname{set}_{d_{g}}([[M]])=\left\{x\in M\,:\,\liminf_{r\to 0}\operatorname{vol}_{g}(B_{d_{g}}(x,r))/r^{m}>0\right\}=M.$
The following lemma depicted in Figure 4.1 is useful for constructing examples:
Lemma 4.11.
Suppose the sphere ${\mathbb{S}}^{2}$ is endowed with the standard round metric, $g_{{\mathbb{S}}^{2}}$, and the singular metric, $g_{h}$,
(4.38) $g_{h}=dr^{2}+\sin^{2}(r)\,d\theta^{2}+(h^{\prime}(r))^{2}dr^{2}\geq g_{{\mathbb{S}}^{2}}=dr^{2}+\sin^{2}(r)\,d\theta^{2}$
defined on a sphere parametrized by $r\in[0,\pi]$ and $\theta\in[0,2\pi]$, where $h(r(x))$ is smooth on $r^{-1}(0,\pi)$, increasing on $[0,\pi/2]$ and decreasing on $[\pi/2,\pi]$ with $h(r)=h(\pi-r)$, but only continuous at the poles $p_{0},p_{\pi}$ where $r=0,\pi$ respectively. We can define a unique compact metric space, $({\mathbb{S}}^{2},d_{h})$, using $g_{h}$-lengths of curves, $C$, such that
(4.39) $C:(0,1)\to{\mathbb{S}}^{2}\setminus\{p_{0},p_{\pi}\}$
with bounded radial arclength functions from the poles $p=p_{0},p_{\pi}$,
[…]
So that $({\mathbb{S}}^{2},d_{h})$ is the metric completion of $({\mathbb{S}}^{2}\setminus\{p_{0},p_{\pi}\},d_{g_{h}})$.
We can find a parametrization of the standard integral current $[[{\mathbb{S}}^{2}]]$ on $({\mathbb{S}}^{2},d_{{\mathbb{S}}^{2}})$ by a countable collection of Lipschitz charts into $({\mathbb{S}}^{2},d_{h})$ which map onto $r^{-1}(0,\pi)$ so that we have an integral current space
(4.41) $\left(\operatorname{set}_{d_{h}}([[{\mathbb{S}}^{2}]]),d_{h},T_{h}\right)\textrm{ where }r^{-1}(0,\pi)\subset\operatorname{set}_{d_{h}}(T_{h})\subset{\mathbb{S}}^{2},$
and where $T_{h}=[[{\mathbb{S}}^{2}]]$ in the sense of Definition 4.9. So $\operatorname{spt}_{d_{h}}(T_{h})={\mathbb{S}}^{2}$ and $\partial T_{h}=0$. The poles are included in $\operatorname{set}_{d_{h}}(T_{h})$ iff
(4.42) $\lim_{r\to 0}|h^{\prime}(r)|<\infty.$
Figure 4.1. Lemma 4.11 applied with $h=h_{cone}$ of Example 4.12 on the left (which will include the singular poles) and with $h=h_{cusp}$ of Example 4.13 on the right (which will not include the singular poles due to the cusp).
Proof.
We easily see that we can define $d_{h}$ away from the poles using lengths of curves that avoid the poles. Near $p_{0}$, we have radial arclengths,
[…]
by the properties of $h$. So
(4.44) $\lim_{r\to 0}s_{p_{0}}(r)=\lim_{r\to 0}r+|h(r)-h(0)|=0$
and similarly for $p_{\pi}$. This allows us to define $({\mathbb{S}}^{2},d_{h})$ as in (4.39) so that the identity map to the standard round sphere is a homeomorphism and so $({\mathbb{S}}^{2},d_{{\mathbb{S}}^{2}})$ is compact with bounded diameter.
The usual integral current, $[[{\mathbb{S}}^{2}]]$, on the standard round sphere $({\mathbb{S}}^{2},d_{{\mathbb{S}}^{2}})$ can be defined using the standard single $(r,\theta)$ chart,
(4.45) $\psi:[0,\pi]\times[0,2\pi]\to{\mathbb{S}}^{2},$
with multiplicity one, but this may not be a Lipschitz chart into $({\mathbb{S}}^{2},d_{h})$ due to the singularities at the poles.
Instead we define a new parametrization with a countable collection of multiplicity one charts avoiding the poles,
(4.46) $\psi_{i}:U_{i}\to{\mathbb{S}}^{2}\textrm{ where }\psi_{i}(r,\theta)=\psi(r,\theta),$
and where $U_{i}$ are defined inductively by
(4.47) $U_{1}=[\delta_{1},\pi-\delta_{1}]\times[0,2\pi)\textrm{ and }\,U_{i+1}=[\delta_{i+1},\pi-\delta_{i+1}]\times[0,2\pi)\setminus U_{i}$
where $\delta_{i}$ decrease to 00 so that these are Lipschitz charts in $({\mathbb{S}}^{2},d_{h})$. Then
(4.48) $T_{h}=\sum_{i=1}^{\infty}\psi_{i\#}[[U_{i}]]$
is a rectifiable current on $({\mathbb{S}}^{2},d_{h})$ because
(4.49) $\sum_{i=1}^{\infty}{\mathcal{H}}^{2}(\psi_{i}(U_{i}))\leq\operatorname{vol}_{g_{h}}({\mathbb{S}}^{2})<\infty.$
We see that $T_{h}=[[{\mathbb{S}}^{2}]]$ in the sense of Definition 4.9 because
(4.50) $T_{h}(\pi_{0},\pi_{1},\pi_{2})=[[{\mathbb{S}}^{2}]](\pi_{0},\pi_{1},\pi_{2})$
for any tuple in $({\mathbb{S}}^{2},d_{{\mathbb{S}}^{2}})$, since missing two points does not affect the integration of tuples. So $(\operatorname{set}_{d_{h}}(T_{h}),d_{h},T_{h})$, is an integral current space. Furthermore
(4.51) $\partial T_{h}=\partial[[{\mathbb{S}}^{2}]]=0.$
In order to determine which points lie in $\operatorname{set}_{d_{h}}(T_{h})$ we use (4.21) and the fact that the mass measure is the volume measure as stated above (4.36). It is easy to see that points where $g_{h}$ is smooth lie in the set, so we conclude (4.41) holds. Taking either pole $p\in\{p_{0},p_{\pi}\}$ and $r=r_{p}$, and applying l’Hopital’s rule twice, we have
[…]
(4.53) $=$ $\lim_{s\to 0}\pi r^{\prime}(s)=\lim_{s\to 0}\pi(1+(h^{\prime}(r))^{2})^{-1/2}.$
So $p\in\operatorname{set}_{d_{0}}(T_{h})$ iff this limit is $>0$ which happens iff (4.42) holds. ∎
Example 4.12.
If we consider $g_{h}$ of Lemma 4.11 defined using
(4.54) $h(r)=h_{cone}(r)=1-{|\sin(r)|}$
then $({\mathbb{S}}^{2},d_{cone})$ is a compact metric space with conical singularities, and $(\operatorname{set}_{d_{cone}}([[{\mathbb{S}}^{2}]],d_{h},[[{\mathbb{S}}^{2}]])$ is an integral current space with
(4.55) $\operatorname{set}([[{\mathbb{S}}^{2}]])={\mathbb{S}}^{2}.$
as depicted in Figure 4.1 because
(4.56) $\lim_{r\to 0}|h^{\prime}(r)|=\lim_{r\to 0}\cos(r)=1<\infty.$
Example 4.13.
If we consider $g_{cusp}=g_{h}$ of Lemma 4.11 defined using
(4.57) $h(r)=h_{cusp}(r)=1-\sqrt{\sin(r)}=1-(\sin^{2}(r))^{1/4}$
then $({\mathbb{S}}^{2},d_{cusp})$ is a compact metric space with cusp singularities at the poles, and the triple $(\operatorname{set}_{d_{cusp}}([[{\mathbb{S}}^{2}]],d_{h},[[{\mathbb{S}}^{2}]])$ is an integral current space with
(4.58) $\operatorname{set}_{d_{cusp}}([[{\mathbb{S}}^{2}]])={\mathbb{S}}^{2}\setminus\{p_{0},p_{\pi}\}.$
as depicted in Figure 4.1 because
(4.59) $\lim_{r\to 0}|h^{\prime}(r)|=\lim_{r\to 0}\frac{1}{2}\frac{\cos(r)}{\sqrt{\sin(r)}}=\infty.$
Theorem 1.1 is careful with the description of the SWIF limit because the space, $(M,d_{\infty},[[M]])$, might not be an integral current space, as it might fail to have $\operatorname{set}([[M]])=M$ as in Example 4.14 depicted in Figure 4.2.
Example 4.14.
Let us consider the sphere, ${\mathbb{S}}^{2}$ with a sequence of metric tensors, $g_{j}=g_{h_{j}}$ as in Lemma 4.11 defined using
(4.60) $h_{j}(r)=1+\tfrac{1}{j}-\left(\tfrac{1}{j^{4}}+\sin^{2}(r)\right)^{1/4}.$
Note that $h_{j}\circ r:{\mathbb{S}}^{2}\to[0,1]$ are smooth functions because $sin^{2}(r)$ is a smooth nonnegative function on a sphere even at the poles where $r=0,\pi$ and $u^{1/4}$ is smooth for $u>0$. In fact
(4.61) $|h_{j}^{\prime}(r)|=\tfrac{1}{4}\left(\tfrac{1}{j^{4}}+\sin^{2}(r)\right)^{-3/4}\left(2\sin(r)\cos(r)\right).$
so
(4.62) $|\nabla(h_{j}\circ r)|=|h_{j}^{\prime}(r)|\leq|h_{j+1}^{\prime}(r)|\leq|h_{cusp}^{\prime}(r)|.$
Thus we have a monotone increasing sequence of metric tensors as described in Remark 2.4. We have uniformly bounded radial arclength parameters. The radial arclengths,
[…]
are uniformly bounded, so we have a uniform upper bound on diameter.
Observe that $h_{j}\to h_{cusp}$ of Example 4.13 smoothly away from the poles. So $d_{j}\to d_{cusp}$ pointwise away from the poles. Since the arclength parameters also converge, we see that $d_{j}\to d_{cusp}$ on ${\mathbb{S}}^{2}\times{\mathbb{S}}^{2}$. Since $({\mathbb{S}}^{2},d_{{\mathbb{S}}^{2}})$ is compact we have all the hypotheses of Theorem 1.1. This is an example where the SWIF limit is a proper subset of the GH limit:
(4.64) $\operatorname{set}_{d_{cusp}}\left([[{\mathbb{S}}^{2}]]\right)\neq{\mathbb{S}}^{2}.$
Figure 4.2. Example 4.2 has $d_{j}\leq d_{j+1}$ because $|h_{j}^{\prime}|\leq|h_{j+1}^{\prime}|$ and converges to Example 4.13 which has the poles with cusp singularities removed.
4.6. Intrinsic Flat Distance
In [SW11], Sormani and Wenger define the intrinsic flat distance between two integral current spaces as follows:
Definition 4.15.
(4.65) $d_{SWIF}((X_{a},d_{a},S_{a}),(X_{b},d_{b},S_{b}))=\inf\{d_{F}^{Z}\left(f_{a\#}(S_{a}),f_{b\#}(S_{b})\right)\}$
where the infimum is taken over all complete metric spaces, $(Z,d_{Z})$, and over all distance preserving maps,
(4.66) $f_{a}:(X_{a},d_{a})\to(Z,d_{Z})\textrm{ and }f_{b}:(X_{b},d_{b})\to(Z,d_{Z}).$
Here $d_{F}^{Z}$ denotes the Flat distance between integral currents $T_{a},T_{b}$ on $(Z,d_{Z})$:
(4.67) $d_{F}^{Z}\left(T_{a},T_{b}\right)=\inf\{{\mathbf{M}}(A)+{\mathbf{M}}(B):A+\partial B=T_{a}-T_{b}\}$
where the infimum is taken over all integral currents $A,B$ on $(Z,d_{Z})$.
We say that a sequence of integral current spaces converges in the intrinsic flat sense,
(4.68) $(X_{j},d_{j},S_{j})\stackrel{{\scriptstyle\textrm{SWIF}}}{{\longrightarrow}}(X_{\infty},d_{\infty},S_{\infty}),$
if
(4.69) $d_{SWIF}((X_{j},d_{j},S_{j}),(X_{\infty},d_{\infty},S_{\infty}))\to 0.$
In [SW11], Sormani-Wenger prove this implies there is a common complete metric space, $(Z,d_{Z})$, and distance preserving maps, $f_{j}:X_{j}\to Z$ such that, $T_{j}=f_{j\#}S_{j}$ converge weakly to $T_{\infty}=f_{\infty\#}S_{\infty}$ as integral currents on $Z$. Thus by Ambrosio-Kirchheim semicontinuity of mass,
(4.70) $\liminf_{j\to\infty}{\mathbf{M}}_{d_{j}}(S_{j})=\liminf_{j\to\infty}{\mathbf{M}}_{d_{Z}}(T_{j})\geq{\mathbf{M}}_{d_{Z}}(T_{\infty})={\mathbf{M}}_{d_{\infty}}(S_{\infty}).$
We say the sequence of integral current spaces converges in the volume preserving intrinsic flat sense if we have (4.68) and
(4.71) ${\mathbf{M}}_{d_{j}}(S_{j})\to{\mathbf{M}}_{d_{\infty}}(S_{\infty}).$
This notion was first studied by Portegies in [Por15]. Additional work was completed by Jauregui-Lee in [JL21]. The consequences of $\mathcal{VF}$ convergence are most recently reviewed by Jauregui-Perales-Portegies in [JPP22].
4.7. Intrinsic Flat Convergence Theorem
We can now state our SWIF convergence theorem:
Theorem 4.16.
Suppose $(X,d_{j})$ is a monotone increasing sequence of metric spaces satisfying all the hypotheses of Theorem 3.2 so that $d_{j}\to d_{\infty}$ uniformly and
(4.72) $(X,d_{j})\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}(X,d_{\infty})\textrm{ which is compact}.$
If there is a current, $T$, which is an integral current structure for $(X,d_{j})$,
(4.73) $\operatorname{set}_{d_{j}}(T)=X\quad\forall j\in\mathbb{N},$
with uniform upper bounds on total mass,
(4.74) ${\mathbf{M}}_{d_{j}}(T)\leq V_{0}\textrm{ and }{\mathbf{M}}_{d_{j}}(\partial T)\leq A_{0}\quad\forall j\in\mathbb{N},$
then
(4.75) $(X,d_{j},T)\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}(M_{\infty},d_{\infty},T_{\infty}).$
where $T_{\infty}=T$ in the sense defined in Definition 4.9 and
(4.76) $M_{\infty}=\operatorname{set}_{d_{\infty}}(T_{\infty})\subset X\textrm{ with }\overline{M}_{\infty}=X.$
Furthermore,
(4.77) ${\mathbf{M}}_{d_{\infty}}(T_{\infty})=\lim_{j\to\infty}{\mathbf{M}}_{d_{j}}(T).$
so we have volume preserving SWIF convergence.
Recall that in Example 4.14 we saw that the SWIF limit, $M_{\infty}$, might be a proper subset of the Gromov-Hausdorff limit, $X$.
4.8. The Proof of Riemannian Theorem 1.1
Before we prove Theorem 4.16, we show in this subsection how it can be applied to prove Theorem 1.1:
Proof.
Note that the hypotheses of Theorem 1.1 imply the hypotheses of Theorem 3.2 and Theorem 4.16 as follows:
  • Monotonicity of metric tensors in (1.1) implies monotonicity of distances as in Definition 2.2 by Lemma 2.3 as in Remark 2.4.
  • Adding the diameter bound in (1.2) implies the pointwise convergence of $d_{j}\to d_{\infty}$ in Lemma 2.5 and the diameter hypothesis of Theorem 3.2.
  • The assumption that $(M,d_{\infty})$ is compact in Theorem 1.1 implies the compactness assumption in Theorem 3.2, so that we have uniform convergence as in Proposition 3.6 and a common metric space $Z_{j}$ for $(M,d_{j})$ and $(M,d_{\infty})$ as in Proposition 3.9 and $(M,d_{j})\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}(M,d_{\infty})$ as in Theorem 3.2.
  • The hypothesis that $(M,g_{j})$ are smooth oriented Riemannian manifolds in Theorem 1.1 implies $(M,d_{g_{j}},[[M]]])$ are integral current spaces with $T=[[M]]$ by (4.37).
  • The volume bounds in (1.4) imply the mass bounds in (4.74) by (4.36).
Thus we may apply the Theorem 4.16 and (4.36) to conclude the volume preserving intrinsic flat convergence in (1.5) to a limit space satisfying (1.6). ∎
4.9. Finding the Current Structure of the SWIF Limit for Theorem 4.16
In [SW11], Sormani-Wenger proved that when one has a sequence of integral current spaces which converge in the Gromov-Hausdorff sense and have uniformly bounded total mass as in (4.74), then SWIF limit of the integral current spaces is a subset of the Gromov-Hausdorff limit. For Theorem 4.16 we wish to show the closure of the SWIF limit is the GH limit, so we need stronger control on the SWIF limit’s current structure. To achieve this, we will repeat the steps of the argument in [SW11] using the special properties of the common metric space, $(Z,d_{Z})$, that we constructed in Proposition 3.9. See Figure 4.3.
Figure 4.3. The distance preserving maps from Proposition 3.9.
Proposition 4.17.
Under the hypotheses Theorem 4.16, we have the metric space $(Z,d_{Z})$ as constructed in Proposition 3.9 depicted in Figure 4.3 with distance preserving maps,
(4.78) $\chi_{j}\,$ $=$ $\zeta_{j}\circ f_{j\,,(j,\infty)}:\,(X,d_{j}\,)\to\,(Z_{j,\infty},d_{Z_{j,\infty}})\to(Z,d_{Z}),$
(4.79) $\chi_{\infty}$ $=$ $\zeta_{j}\circ f_{\infty,(j,\infty)}:(X,d_{\infty})\to(Z_{j,\infty},d_{Z_{j,\infty}})\to(Z,d_{Z}),$
such that the pushforwards
(4.80) $T_{j}=\chi_{j\#}(T)=\zeta_{j\#}(f_{j\#}(T))$
are integral currents in $Z$ whose supports are
(4.81) $\operatorname{spt}(T_{j})=\zeta_{j}(f_{j}(X))=\zeta_{j}(X\times\{h_{j}\})\subset\zeta_{j}(Z_{j,\infty})\subset Z$
and have
(4.82) ${\mathbf{M}}_{d_{Z}}(T_{j})={\mathbf{M}}_{d_{j}}(T)\leq V_{0}\textrm{ and }{\mathbf{M}}_{d_{Z}}(\partial T_{j})={\mathbf{M}}_{d_{j}}(\partial T)\leq A_{0}.$
Furthermore, there is a subsequence $T_{j_{k}}$ which converges weakly as currents in $(Z,d_{Z})$ to an integral current, $T_{\infty}$, whose support lies on the central spine in $Z$,
(4.83) $\operatorname{spt}(T_{\infty})=Z_{0}=\chi_{\infty}(X)=\zeta_{j}(X\times\{0\})\subset\zeta_{j}(Z_{j,\infty})\subset Z.$
Finally $T=\chi_{\infty\#}^{-1}T_{\infty}$ is an integral current on $(X,d_{\infty})$ as in Definition 4.9, and
(4.84) ${\mathbf{M}}_{d_{\infty}}(T)=\lim_{k\to\infty}{\mathbf{M}}_{d_{j_{k}}}(T).$
Note that before we prove this proposition, we do not yet know that $T$ is an integral current on $(X,d_{\infty})$. See Remark 4.8. After we prove this proposition we will apply it to prove the same statement holds without needing a subsequence in Corollary 4.18.
Proof.
Since the maps in (4.78) are distance preserving, the pushforward integral currents $T_{j}$ on $(Z,d_{Z})$ defined as in (4.80), satisfy (4.82) and (4.81) following the definitions of pushforward and mass (cf. [SW11]).
Since $(Z,d_{Z})$ is compact, we can apply the Ambrosio-Kirchheim Compactness Theorem, Theorem 4.10, to conclude that there is a subsequence $T_{j_{k}}$ which converges weakly as currents in $(Z,d_{Z})$ to an integral current, $T_{\infty}$.
Figure 4.4. Here we see integral currents $T_{j}$ in $Z$ and $B_{d_{Z}}(z,t)$ of the proof that $\operatorname{spt}(T_{\infty})\subset Z_{0}=\chi_{\infty}(X)\subset Z$ as claimed in (4.85).
First we claim
(4.85) $\operatorname{spt}(T_{\infty})\subset Z_{0}=\chi_{\infty}(X)=\zeta_{j}(X\times\{0\}\subset Z_{j,\infty})\subset Z.$
Take any $z\in Z\setminus Z_{0}$ as in Figure 4.4. Then there exists $j_{z}\in{\mathbb{N}}$ such that $z=\zeta_{j_{z}}(f_{j_{z}}(x,t))$ where $t>0$. Then by the definition of $d_{Z}$ in Proposition 3.9, the ball:
(4.86) $B_{d_{Z}}(z,t)\subset\zeta_{j_{z}}(f_{j_{z}}(X\times(0,h_{j_{z}}])))\subset\zeta_{j_{z}}(Z_{j_{z}})\setminus Z_{0}.$
In particular
(4.87) $B_{d_{Z}}(z,t)\cap\zeta_{j}(f_{j}(X\times\{h_{j}\}))=\emptyset$
for all $j>j_{z}$. By the lower semicontinuity of mass under weak convergence as in (4.34) we have
(4.88) $||T_{\infty}||(B_{d_{Z}}(z,t))\leq\liminf_{j\to\infty}||T_{j}||(B_{d_{Z}}(z,t))=0$
so $z\notin\operatorname{spt}(T_{\infty})$. Thus $\operatorname{spt}(T)\subset Z_{0}$.
Since $\chi_{\infty}:(X,d_{\infty})\to(Z_{0},d_{Z})$ is an isometry, we can define:
(4.89) $S=\chi^{-1}_{\infty\#}T_{\infty}$
as an integral current on $(X,d_{\infty})$.
We claim that $S=T$ in the sense of Definition 4.9. Given any tuple, $(\pi_{0},\pi_{1},...,\pi_{m})$ of Lipschitz functions on $(X,d_{j})$, we need only show
(4.90) $T(\pi_{0},\pi_{1},...,\pi_{m})=S(\pi_{0},\pi_{1},...,\pi_{m}).$
By Lemma 4.1 these $\pi_{i}$ are also tuples of Lipschitz functions on $(X,d_{\infty})$, so $S(\pi_{0},...,\pi_{m})$ is well defined.
Figure 4.5. Here we see $W_{j}\subset Z$ of Lemma 3.10 in dark gray, the rest of $Z$ is lighter, and $Z_{0}\subset W_{j}\subset Z$ is in black with the isometry, $\chi^{-1}_{\infty}$, to $(X,d_{\infty})$.
Recall the set $W_{j}\subset Z$ from Lemma 3.10, Lemma 4.3, and Remark 4.4 depicted in Figure 4.5. By Lemma 4.3, we define a tuple of Lipschitz functions, $(\tilde{\pi}_{0},\tilde{\pi}_{1},...,\tilde{\pi}_{m})$, on the subset, $(W_{j},d_{Z})$, inside $(Z,d_{Z})$, such that
(4.91) $\tilde{\pi}_{i}(\zeta_{j_{k}}(x,t))=\pi_{i}(x)\qquad\forall x\in X,\,\,t\in[0,h_{j_{k}}],\,\,j_{k}\geq j.$
By (4.89) and the fact that
(4.92) $\tilde{\pi}_{i}(\zeta_{j_{k}}(x,t))=\pi_{i}(x)=\tilde{\pi}_{i}(\chi_{\infty}(x)))\qquad\forall x\in X,\,\,t\in[0,h_{j_{k}}],\,\,j_{k}\geq j,$
we have by definition of pushforward that
(4.93) $S(\pi_{0},\pi_{1},...,\pi_{m})$ $=$ $(\chi^{-1}_{\infty\#}T_{\infty})(\tilde{\pi}_{0}\circ\chi_{\infty},\tilde{\pi}_{1}\circ\chi_{\infty},...,\tilde{\pi}_{m}\circ\chi_{\infty})$
(4.94) $=$ $T_{\infty}(\tilde{\pi}_{0},\tilde{\pi}_{1},...,\tilde{\pi}_{m}).$
Note that for any $j_{k}\geq j$ we have
(4.95) $\operatorname{spt}(T_{j_{k}})\subset\zeta_{j_{k}}(Z_{j_{k}})\subset W_{j}\subset Z.$
So $T_{j_{k}}$ converge weakly as currents to $T_{\infty}$ in $(W_{j},d_{Z})$: for any Lipschitz tuple, $\omega$, on $(W_{j},d_{Z})$, we have
(4.96) $T_{j_{k}}(\omega)\to T_{\infty}(\omega).$
Thus we have
(4.97) $S(\pi_{0},\pi_{1},...,\pi_{m})=\lim_{k\to\infty}T_{j_{k}}(\tilde{\pi}_{0},\tilde{\pi}_{1},...,\tilde{\pi}_{m}).$
On the other hand, by (4.91), we have,
(4.98) $\tilde{\pi}_{i}(\zeta_{j_{k}}(x,t))=\pi_{i}(x)=\tilde{\pi}_{i}(\zeta_{j_{k}}(f_{j_{k}}(x))))\quad\forall x\in X,\,\,t\in[0,h_{j_{k}}],\,\,j_{k}\geq j.$
Thus $T_{j_{k}}=\zeta_{j_{k}\#}f_{j_{k}\#}T$ and we have
(4.99) $T_{j_{k}}(\tilde{\pi}_{0},...,\tilde{\pi}_{m})$ $=$ $(\zeta_{j_{k}\#}f_{j_{k}\#}T)(\tilde{\pi}_{0}\circ\zeta_{j_{k}}\circ f_{j_{k}},...,\tilde{\pi}_{m}\circ\zeta_{j_{k}}\circ f_{j_{k}})$
(4.100) $=$ $T(\pi_{0},...,\pi_{m}).$
Substituting this into (4.97), and taking the limit $j_{k}\to\infty$, we reach our claim in (4.90). We conclude that $S=T$ in the sense of Definition 4.9.
Now we claim $\operatorname{spt}_{d_{Z}}(T_{\infty})=Z_{0}$. Since we know $T_{\infty}=\chi_{\infty\#}T$, this is the same as showing $\operatorname{spt}_{d_{\infty}}(T)=X$. Given any $x\in X$ and any $r>0$ we have
(4.101) $||T||_{d_{\infty}}(B_{d_{\infty}}(x,r))\geq||T||_{d_{j}}(B_{d_{j}}(x,r))$
by Lemma 4.7. By our hypotheses, we have uniform convergence of the distance functions, $|d_{j}-d_{\infty}|<\epsilon_{j}\to 0$, so for all $j$ such that $\epsilon_{j}<r/2$
(4.102) $B_{d_{j}}(x,r)\supset B_{d_{j}}(x,r-\epsilon_{j}).$
Combining this with the above,
(4.103) $||T||_{d_{\infty}}(B_{d_{\infty}}(x,r))\geq||T||_{d_{j}}(B_{d_{j}}(x,r-\epsilon_{j}))>0.$
Thus $\operatorname{spt}_{d_{\infty}}(T)=X$ and $\operatorname{spt}_{d_{Z}}(T_{\infty})=Z_{0}$.
Finally, by Ambrosio-Kircheim’s semicontinuity of mass combined with the fact that $\zeta_{\infty}$ is distance preserving we have
(4.104) ${\mathbf{M}}_{d_{\infty}}(T)={\mathbf{M}}_{d_{Z}}(T_{\infty})\leq\liminf_{k\to\infty}{\mathbf{M}}_{d_{Z}}(T_{j_{k}})=\liminf_{k\to\infty}{\mathbf{M}}_{d_{j_{k}}}(T).$
Yet by Lemma 4.7, we know
(4.105) ${\mathbf{M}}_{d_{\infty}}(T)\geq{\mathbf{M}}_{d_{j_{k}}}(T)$
so we have mass convergence. ∎
Corollary 4.18.
Under the hypotheses of Proposition 4.17, we have the conclusions $T_{j}=\chi_{j\#}T$ converge weakly to $T_{\infty}=\chi_{\infty\#}T$ as integral currents in $(Z,d_{Z})$ and
(4.106) ${\mathbf{M}}_{d_{\infty}}(T)=\lim_{j\to\infty}{\mathbf{M}}_{d_{j}}(T).$
without taking a subsequence.
Proof.
First note that (4.106) follows from (4.84) combined with the monotonicity and Lemma 4.7.
To prove weak convergence, assume on the contrary that there is a tuple, $(\pi_{0},...\pi_{m})$, on $(Z,d_{Z})$, and a $\delta>0$ with a subsequence, we also denote $T_{j}$, such that
(4.107) $|T_{j}(\pi_{0},...\pi_{m})-T_{\infty}(\pi_{0},...\pi_{m})|<\delta\quad\forall j\in\mathbb{N}.$
Note that this subsequence $T_{j}$ satisfies the hypotheses of Proposition 4.17, so it has a subsequence which converges weakly to $T$, contradicting (4.107). ∎
4.10. Proving SWIF Convergence
In this section we complete the proof of Theorem 4.16:
Proof.
By Corollary 4.18 of Proposition 4.17, we have pushforwards, $T_{j}=\chi_{j\#}T$, converging weakly to $T_{\infty}=\chi_{\infty\#}T$ as integral currents in $Z$.
By the uniform upper bound on total mass, the compactness of $(Z,d_{Z})$ and Wenger’s Flat=Weak Convergence theorem in [Wen07] (applied to metric spaces as explained in [SW11]) there exist integral currents, $B_{j},A_{j}$, on $(Z,d_{Z})$ such that
(4.108) $\partial B_{j}+A_{j}=T_{j}-T_{\infty}\textrm{ and }{\mathbf{M}}_{d_{Z}}(B_{j})+{\mathbf{M}}_{d_{Z}}(A_{j})\to 0$
so that
(4.109) $d_{F}^{Z}(\chi_{j\#}T,\chi_{\infty\#}T)\to 0.$
Taking $M_{\infty}=\operatorname{set}_{d_{\infty}}(T)$ and applying the fact that $\chi_{j}$ and $\chi_{\infty}$ are distance preserving maps, we have
(4.110) $d_{SWIF}((X,d_{j},T),(M_{\infty},d_{\infty},T))\to 0.$
Since we have the convergence of masses in Corollary 4.18, we have $\mathcal{VF}$ convergence as well. ∎
Remark 4.19.
It is an exercise to circumvent applying Wenger’s Flat=Weak Convergence Theorem in the above proof, by explicitly constructing a sequence of integral currents, $A_{j}$ and $B_{j}$, on $(Z,d_{Z})$ using the special properties of $(Z,d_{Z})$ proven in Prop 3.9. Since we have shown $T$ is an integral current on $(X,d_{\infty})$, there are integers, $a_{i}$, and charts
(4.111) $\psi_{i}:(U_{i}\subset{\mathbb{R}}^{m},d_{{\mathbb{R}}^{m}})\to(X,d_{\infty})\textrm{ with }\sum_{i=1}^{\infty}|a_{i}|\,\mathcal{H}^{m}_{d_{\infty}}(\psi_{i}(U_{i}))<\infty$
such that
(4.112) $T=\sum_{i=1}^{\infty}a_{i}\,\psi_{i\#}[[U_{i}]],$
and the same holds for $\partial T$. Then it is natural to define the integral current, $B_{j}$, using the same integers $a_{i}$ and charts,
(4.113) $\tilde{\psi}_{i,j}:(U_{i}\times[0,h_{j}]\subset{\mathbb{R}}^{m+1},d_{{\mathbb{R}}^{m+1}})\to(Z,d_{Z})$
such that
(4.114) $\tilde{\psi}_{i,J}(v,t)=\zeta_{j}(\psi_{i}(v),t).$
We then similarly construct the integral current, $A_{j}$, using the charts for $\partial T$ (which are not necessarily the boundaries of charts of $T$). Next we would verify that both atlases of charts are Lipschitz and that $A_{j}$ and $B_{j}$ have finite masses satisfying
(4.115) ${\mathbf{M}}_{d_{Z}}(A_{j})\leq\,C\,h_{j}\,{\mathbf{M}}_{d_{\infty}}(T)\textrm{ and }{\mathbf{M}}_{d_{Z}}(B_{j})\leq\,C\,h_{j}\,{\mathbf{M}}_{d_{\infty}}(T).$
Most difficult is showing $A_{j}+\partial B_{j}=T_{j}-T_{\infty}$ by texting against arbitrary Lipschitz tuples on $(Z,d_{Z})$. It may be necessary to restrict to $W_{j}$ to get everything to work.
5. Open Problems
We know that without the compactness assumption in our Theorem 1.1 and Theorem 4.16, that we may not have Gromov-Hausdorff Convergence and a common compact metric space. However we know that by Wenger’s Compactness Theorem the diameter and mass bounds imply that there is a subsequence converging to a SWIF limit,
(5.1) $(X,d_{j_{k}},T)\stackrel{{\scriptstyle\textrm{SWIF}}}{{\longrightarrow}}(X^{\prime}_{\infty},d^{\prime}_{\infty},T^{\prime}_{\infty}).$
Is it possible that with monotonicity, $d_{j}\leq d_{j+1}$, one can show any of the following:
  • the sequence itself converges,
  • there is volume preserving intrinsic flat convergence,
  • the limit, $X^{\prime}_{\infty}\subset X$, with restricted metric, $d^{\prime}_{\infty}=d_{\infty}$, possibly with closure, $\bar{X}_{\infty}=X$, and maybe $T_{\infty}=T$.
Note that assuming any of the following additional hypotheses might help:
  • The integral current spaces, $(X,d_{j},T)$, are Riemannian, $(M,d_{g_{j}},[[M]])$, possibly without boundary.
  • The integral current spaces, $(X,d_{j},T)$, have empty boundary, $\partial T=0$, so we don’t have to use $A_{j}$ in the proof of SWIF convergence.
  • The integral current structure, $T$, is assumed to be an integral current on $(X,d_{\infty})$ as an hypothesis, and that perhaps also that it has a bound on its multiplicity.
It would be interesting to find counter examples to any or all of these statements as well.